Use the system of equations to solve the following:
The sum of two numbers is 73. When the smaller number is subtracted from twice the greater number the result is 50. Find the numbers.

Answer by mananth(16949)

let smaller number be x
the other number be y
x+y=73.........................1
2y-x=50........................2
Add the equations
3y=123
/3
y = 123/3
y=41..
plug value of y in (1)
x+41=73
-41
x=73-41
x=32
...
x=32, y =41
...
m.ananth@hotmail.ca